2. A pure ALOHA network transmits 400-bit frames on a shared channel of 200 kbps. What is the throughput if the system (all stations together) produces

a. 1000 frames per second

b. 500 frames per second

c. 250 frames per second
Answer #1

Throughput of PURE ALOHA is S = G X e-2G  

Here G is average number of frames generated during one frame transmission time.

Given, from question framesize = 400 bits

bandwidth of channel = 200 kbps

Transmission time =  frame-size / bandwidth

= 400 / 200 ms

= 2 ms

a. if system creates 1000 frames per second then it creates 1 frame per milli-second.

this implies G=1/2 as transmission time is 2ms

S= G x e-2G = 0.184

this means throughput = 1000 x 0.184 =184 frames

only 184 frames out of 1000 will reach probably at receiver.

b. if system creates 500 frames per second then it creates 1/2 frames per milli-second.

this implies G=1/4 as transmission time = 2ms

S= G x e-2G = 0.152

this means throughput = 500 x 0.152 = 76 frames

only 76 frames out of 500 will reach probably at receiver.

c. if system creates 250 frames per second then it creates 1/4 frames per milli second.

this implies G=1/8 as transmission time=2ms

S= G x e-2G = 0.110

this means throughput = 250 x 0.110 = 27 frames

only 27 frames out of 250 will reach probably at receiver end.